A twisted steel wire of length L,cross sectional area A and modulus of elasticity E in tension . The wire is stretched horizontally between two fixed points A and B but without any initial tension . A load P is then suspended from the mid -pointC of the wire .find the vertical deflection of the mid-point C of the wire . assume that the deflection is small compared to the length of the wire.


1
Expert's answer
2021-01-19T02:32:18-0500

Workdone ="\frac{1}{2}" ​× Stress × Strain × Volume 

Workdone ="\frac{1}{2}" ​×Y×(strain)2×Volume

"Workdone = 
\frac{1}{2} ×Y×( \frac{x}{l} ) 2×AL"